OpenStudy (anonymous):

What is the center and the radius of the circle given? x^2+y^2-12x-6y+9=0

OpenStudy (anonymous):

So far I got x^2+y^2-12x-6y+9=0 x^2-12x+y^2-6y+9=0 (x^2-12x+36)+(y^2-6y+9)+9-36-9=0 I don't know what to do after that

OpenStudy (anonymous):

http://prntscr.com/2lmpz2

OpenStudy (anonymous):

I don't get how that is suppose to help me?

OpenStudy (anonymous):

Nevermind I got it thanks!